OpenStudy (anonymous):

simplify by taking the roots of the denominator and numerator, assume that all the expressions under radicals represent positive numbers. 3sqrt 64x^14/y^3

OpenStudy (anonymous):

Im not understanding this question.

OpenStudy (anonymous):

Is the question\[\frac{3 \sqrt{64x^{14}}}{y^3}\]

OpenStudy (anonymous):

Or maybe\[\sqrt[3]{\frac{64x^{14}}{y^3}}\]

OpenStudy (anonymous):

its the second one.

OpenStudy (anonymous):

OK, here we go. Factor first, breaking things down into powers with multiples of three where we can.\[\sqrt[3]{\frac{64x^{14}}{y^3}}=\sqrt[3]{\frac{2^6x^{12}x^2}{y^3}}\]See that?

OpenStudy (anonymous):

Then take powers of three outside the radical. \[\sqrt[3]{\frac{2^6x^{12}x^2}{y^3}}=\frac{\sqrt[3]{2^6x^{12}x^2}}{\sqrt[3]{y^3}}=\frac{2^3x^4 \sqrt{x^2}}{y}=\frac{8x^4\sqrt[3]{x^2}}{y}\]

OpenStudy (anonymous):

"Powers of three" is probably a bad way to say that. Powers with exponent factors of three is more like it.

OpenStudy (anonymous):

oh i see it, I worked it out, but I didnt get that answer, so im glad I asked.

OpenStudy (anonymous):

Do math every day.

OpenStudy (anonymous):

you do math everyday? Yea im taking Alg.2B for college, and its kicking my butt